OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)?

OpenStudy (anonymous):

Hint: (f+g) 5 = f(5) + g(5)

OpenStudy (anonymous):

(f+g)(5)=13*25+35-5=355